What are the root or imaginary numbers of: (2x-5)^2=25
What I did: 2x-5 = plus or minus the square root of 25
then my final answer was : x = (5 plus or minus the square root of 25)/2
sorry i do not have a plus or minus key nor do i have a square root key

Answer by jim_thompson5910(35256)

The square root of 25 is 5.

So you will have

x = (5 plus or minus 5)/2

which effectively breaks into these two equations

x = (5+5)/2 or x = (5-5)/2

x = 10/2 or x = 0/2

x = 5 or x = 0

Therefore, the two solutions are x = 5 or x = 0